OpenStudy (eric_d):

A stone is projected vertically upward with a velocity of 10ms^-1 near the edge of the cliff.Given height of cliff=20mTaking g as 10ms^-2,find a)time taken for the stone to reach the ground b)velocity of stone when it hits the ground c)sketch the s-t,v-t and a-t graphs

OpenStudy (eric_d):

a)3.2s b)-22ms^-1

OpenStudy (eric_d):

c)sketch the s-t,v-t and a-t graphs

ganeshie8 (ganeshie8):

3.2 is correct for part a

OpenStudy (eric_d):

Yes.. @abhisar taught me 2 days back I'm stuck with c)

ganeshie8 (ganeshie8):

-22 ms^-1 is also correct for part b

ganeshie8 (ganeshie8):

so your equation for height of stone above ground is : \[\large s(t) = 20 + 10t + \dfrac{1}{2}(-10)t^2\] right ?

OpenStudy (eric_d):

yes

ganeshie8 (ganeshie8):

which is same as : \[\large s(t) = -5t^2 + 10t + 20\]

ganeshie8 (ganeshie8):

doesn't that look like a familiar parabola facing downwards ?

OpenStudy (eric_d):

ys

ganeshie8 (ganeshie8):

so basically your \(s-t\) graph is just a parabola, can you sketch it ?

OpenStudy (eric_d):

will try..

ganeshie8 (ganeshie8):

find its vertex first, so start by changing the equation into vertex form :y = a(x-h)^2+k

OpenStudy (eric_d):

I need to use completing to square to change it to that form

OpenStudy (eric_d):

the*

ganeshie8 (ganeshie8):

yep !

ganeshie8 (ganeshie8):

\(\large s(t) = -5t^2 + 10t + 20\) \(\large ~~~~~~ = -5(t^2 - 2t) + 20\) \(\large ~~~~~~ = -5(t^2 - 2(1)t + \color{Red}{1^2}) + 20 - (\color{red}{-5*1^2})\) \(\large ~~~~~~ = -5(t-1)^2 + 20 +\color{red}{5}\) \(\large ~~~~~~ = -5(t-1)^2 + 25\)

ganeshie8 (ganeshie8):

vertex = (1, 25) find the x/t intercepts and sketch it

ganeshie8 (ganeshie8):

if you want, you may also find y intercept as well to be a bit more precice

OpenStudy (eric_d):

ok..

OpenStudy (abhisar):

Will it be parabola facing upward or downwards ?

OpenStudy (eric_d):

downward

OpenStudy (eric_d):

 this's the answer frm the book